# A postgraduate laboratory experiment to set up a single-photon detector using MKIDs

### v2

@@ 4.2 Development of the data acquisition system
Another experimental activity that was carried out alongside the MKIDs characterization involved developing the data acquisition (DAQ) system for the custom photon detection setup.
As a first step, the FSW-0010 synthesizers were set to the MKIDs resonant frequencies, which fall within the $5-6$ GHz range, in order to generate the probe signals.
The PXIe-5170R digitizer by National Instruments, which can sample frequencies up to $250$ MSample/s/ch, was used for data acquisition. All four channels ($I_{1}$, $Q_{1}$, $I_{2}$, $Q_{2}$) were used to exploit the multiplexing system. To determine the acquisition parameters such as the trigger, number of records to store, record length, and sample rate, a detailed signal analysis is required. Typical values used for data acquisition are $5\times 10^{7}$ Hz for the sampling frequency, $6000$ points for record length, and $1000$ records per acquisition. The digitizer ADCs (Analog to Digital Converters) were configured to work within the range of ±$\pm 5$ V and amplitude resolution of $14$ bits.
One of the most critical aspects of data acquisition is the trigger, which should efficiently select events while rejecting background noise or irrelevant processes. The most significant issue for the trigger used in this experiment was time jitter, which results in signals with the same amplitude triggering acquisition at different times. In order to correctly align the signals, which is crucial for the following data analysis, a moving average smoothing technique was first applied to reduce noise, as shown in fig. 6(a).
@@
The Savitzky-Golay filter [19] was then employed to smooth the second derivative of the signal. The filter works by fitting a polynomial function to a moving window of adjacent data points and substituting the central point of the window with the fitted value. It is then possible to compute the signal n-th derivative from the polynomial fit.
@@ 4.3 IQ correction
The goal of this activity is to transform the raw acquired MKIDs resonances to correctly rotated and centered circles in the IQ plane. This correction will then be applied to each signal observed by the photon detectors. The following steps were employed:
@@
both synthesizers and mixers are subject to a frequency-dependent offset. To address this issue, the RF port was disconnected, and the acquisition was performed for different LO frequencies. The resulting profile was subtracted from the signal of interest.
@@
Given sinusoidal LO and RF signals with similar frequencies, the response of an ideal IQ mixer consists of a circular pattern in the IQ plane. In a realistic IQ mixer, however, the phase between I and Q is not precisely $90^{\circ}$ and the output is not balanced between the two ports. This results in observing an ellipse in the IQ plane. To calibrate each IQ mixer, two synthesizers at slightly different frequencies were connected to the LO and RF ports. The resulting signal was fitted with an ellipse, which was then mapped to a circle.
@@
even far from the MKIDs resonances, signals produced by the synthesizer arrive at the RF port with different amplitudes and phases. This leads to a frequency-dependent background which can be removed by fitting it with polynomial functions in a wide-frequency scan around the resonance. The fitted amplitude $A(\omega)$ and phase $\phi(\omega)$ are then used to correct the measured $S_{21}$ for the on-resonance frequencies:
@@
: each resonance is fitted with a circle, which is then rotated around its center in order to align the resonance’s gap with the I-axis.
@@
the asymmetry in the shape of the resonance appears as a rotation of its circle in the IQ plane [13]. To correct this effect, a counter-rotation and a contraction must be applied:
@@
4.4 Single photon analysis
@@
A more precise energy estimator than the simple peak-to-peak height can be obtained by applying the Optimum Filter (OF) method [7]. The OF is able to achieve the best signal-to-noise ratio under the assumption that the signal is a linear combination of a mean signal and ergodic noise. This method allows the construction of a frequency filter that suppresses the frequencies where the noise contribution is more significant. There are various implementations of the filter, with the one used in this case being based on the Discrete Fourier Transform (DFT).
@@
The expected energy spectrum is shown in fig. 10. When sending $N$ photons to a detector, the number of photons that reach it follows a Poisson distribution with mean $\mu=N$. Assuming a Gaussian energy spread due to the finite detector’s resolution, the observed spectrum is given by a sum of Gaussians weighted with a Poissonian distribution.
@@
$\textrm{model}(\textrm{OFF},\sigma,\mu,A,\textrm{shift},E_{\gamma})=A\sum_{n=0}^{N_{max}}P(n;\mu)\cdot G(\textrm{OFF};nE_{\gamma}+\textrm{shift};\sigma)$ (9)
where $A$ is a normalization factor, $N_{max}$ is a fixed maximum number of considered photons, shift is a global shift of the fit function from zero, $G$ is the Gaussian distribution, $P$ is the Poisson distribution and $E_{\gamma}$ is the conversion from A.U. to eV that can be obtained knowing the energy of one IR photon, i.e. $0.8$ eV.
The measured distribution is shown in fig. 11. The peaks due to different photon numbers cannot be separated, likely due to the detector resolution being much larger than the photons energy.
@@
The average number of photons can still be estimated with the fit. Due to the large number of parameters, it can be difficult to make the fit converge and more than one combination of parameters can give good $\chi^{2}$ values. To avoid that, some parameters can be constrained. For example, $\sigma$ is set to the detector’s resolution, which is calculated as the Root Mean Squared (RMS) of the noise signals (after the OF) since it is the dominant component. All constraints are shown in the table in fig. 11.
The fit is conducted with iminuit [5], which is a Python implementation of the Minuit2 C++ library [11]. The result of the fit, with a $\chi^{2}/$dof$=1.2$, is shown in fig. 11. The best estimate for the average number of photons is $12\pm 2$.
@@ 5 Conclusions
This paper presented a laboratory experiment designed for postgraduate students in the field of quantum technologies, which aims to detect single photons in the far infrared range using Microwave Kinetic Inductance Detectors (MKID). The experiment involves characterizing the resonators and developing a data acquisition system to acquire the signals, concluding with the data analysis of the measurements using the characterized MKIDs.
The experience was first realized at the Università degli Studi di Milano-Bicocca in 2022 and the students reported an average number of observed photons of $12\pm 2$, even if the peaks due to different numbers of photons could not be resolved. Improvements in the experimental setup and the fabrication process of the MKIDs could enhance the resolution, allowing the distinction of different numbers of interacting photons as separate peaks in the energy spectrum.
@@ Acknowledgments
A special thanks goes to our laboratory supervisors for the opportunity and their guidance in our experimental activities. We acknowledge Benno Margergesin and the FBK cleanroom team for the chip fabrication. Additionally we also acknowledge Renato Mezzena and Andrea Vinante for providing the device.
@@
Mention of commercial products is for information only and does not imply recommendation or endorsement.

### v3

@@ 4.2 Development of the data acquisition system
Another experimental activity that was carried out alongside the MKIDs characterization involved developing the DAQ system for the custom photon detection setup.
To generate the probe signals, the two FSW-0010 synthesizers were set to the resonant frequencies of the MKIDs with the highest $Q_{tot}$.
The PXIe-5170R digitizer by National Instruments, which can reach sampling frequencies up to $250$ MHz per channel, was used for data acquisition, employing all four channels ($I_{1}$, $Q_{1}$, $I_{2}$, $Q_{2}$) to exploit the multiplexing system. An analysis of the signals was performed to determine the acquisition parameters such as the trigger, number of records to store, record length, and sample rate. Typical values used for data acquisition are $50$ MHz for the sampling frequency, $6000$ points for record length, and $1000$ records per acquisition. The digitizer ADCs (Analog to Digital Converters) were configured to work within the range of ±$\pm 5$ V and amplitude resolution of $14$ bits.
One of the most critical aspects of data acquisition is the trigger, which should efficiently select events while rejecting background noise or irrelevant processes. The most significant issue for the trigger used in this experiment was time jitter, which results in signals with the same amplitude triggering acquisition at different times. To correctly align the signals, which is crucial for the following data analysis, a moving average smoothing technique was first applied to reduce noise, as shown in fig. 7(a).
@@
The Savitzky-Golay filter [SavGol] was then employed to smooth the second derivative of the signal. The filter works by fitting a polynomial function to a moving window of adjacent data points and substituting the central point of the window with the fitted value. It is then possible to compute the signal n-th derivative from the polynomial fit.
@@ 4.3 IQ correction
The goal of this activity is to write the Python code required for transforming the raw acquired MKIDs resonances to correctly rotated and centered circles in the IQ plane. This correction will then be applied to each signal observed by the photon detectors. The following steps are employed:
@@
both synthesizers and mixers are subject to a frequency-dependent offset. To address this issue, the RF port is disconnected, and the acquisition is performed for different LO frequencies. The resulting profile is subtracted from the signal of interest.
@@
Given sinusoidal Local Oscillator (LO) and RF signals with similar frequencies, the response of an ideal IQ mixer consists of a circular pattern in the IQ plane. In a realistic IQ mixer, however, the phase between I and Q is not precisely $90^{\circ}$ and the output is not balanced between the two ports. This results in observing an ellipse in the IQ plane. To calibrate each IQ mixer, two synthesizers at slightly different frequencies are connected to the LO and RF ports. The resulting signal is fitted with an ellipse, which is then mapped to a circle.
@@
even far from the MKIDs resonances, signals produced by the synthesizer arrive at the RF port with different amplitudes and phases. This leads to a frequency-dependent background which can be removed by fitting a wide-frequency scan around the resonance with the model of eq. 7, approximating the background as a polynomial. The fitted amplitude $A(\omega)$ and phase $\phi(\omega)$ are then used to correct the measured $S_{21}$ for the on-resonance frequencies:
@@
: each resonance is fitted with a circle, which is then rotated around its center to align the resonance’s gap with the I-axis.
@@
the asymmetry in the shape of the resonance appears as a rotation of its circle in the IQ plane [khalil_analysis_2012]. To correct this effect, a counter-rotation and a contraction must be applied:
@@
4.4 Photon-counting analysis
@@
A more precise energy estimator than the simple peak-to-peak height can be obtained by applying the Optimum Filter (OF) method [Gatti1986]. The OF is able to achieve the best signal-to-noise ratio under the assumption that the signal is a linear combination of a mean signal and ergodic noise. This method allows the construction of a frequency filter that suppresses the frequencies where the noise contribution is more significant. There are various implementations of the filter, with the one used in this case being based on the Discrete Fourier Transform (DFT).
@@
The expected energy spectrum is shown in fig. 10. The laser light reaching the detector consists of a coherent state given by a superposition of different photon-number states, which are distributed according to Poisson statistics with average $\mu$. Since each event observed by the detector arises from a definite number of photons being absorbed by the resonator, the spectrum reconstructed by an ideal detector with infinite energy resolution reflects the structure of the coherent state, consisting of a series of discrete peaks at integer multiples of the laser’s single-photon energy.
Assuming a Gaussian energy spread due to the finite detector’s resolution, the observed spectrum is given by a sum of Gaussians weighted by a Poissonian distribution.
@@
$\textrm{model}(\textrm{OFF},\sigma,\mu,A,\delta,E_{\gamma})=A\sum_{n=0}^{N_{max}}P(n,\mu)\cdot G(\textrm{OFF},nE_{\gamma}+\delta,\sigma)$ (9)
where $A$ is a normalization factor, $N_{max}$ is a fixed maximum number of considered photons, $\delta$ is a global shift of the fit function from zero, $G$ is the Gaussian distribution, $P$ is the Poisson distribution and $E_{\gamma}$ is the conversion from A.U. to eV that can be obtained knowing the energy of one infrared photon, i.e. $0.8$ eV.
The measured distribution is shown in fig. 11. The peaks due to different photon numbers cannot be separated, likely due to the detector resolution being much larger than the photon energy.
@@
The average number of photons can still be estimated with the fit. Due to the large number of parameters, it can be difficult to make the fit converge and more than one combination of parameters can give good $\chi^{2}$ values. To avoid that, some parameters can be constrained. For example, $\sigma$ is set to the detector’s resolution, which is calculated as the Root Mean Squared (RMS) of the noise signals (after the OF) since it is the dominant component. All constraints are shown in the table in fig. 11. The result of the fit, with a $\chi^{2}/$dof$=1.2$, is shown in fig. 11. The best estimate for the average number of photons is $12\pm 2$.
@@ 5 Conclusions
This paper presented a laboratory experiment designed for graduate students in the field of quantum technologies, which aims to measure the number-resolved energy spectrum of photons in the far-infrared range using Microwave Kinetic Inductance Detectors (MKID). The experiment involves characterizing the resonators and developing a data acquisition system to acquire the signals, concluding with the data analysis of the measurements using the characterized MKIDs.
The experience was first realized at the University of Milano-Bicocca in 2022 and the students reported an average number of observed photons of $12\pm 2$, even if the peaks due to different numbers of photons could not be resolved. Improvements in the experimental setup and the fabrication process of the MKIDs could enhance the resolution, allowing the distinction of different numbers of interacting photons as separate peaks in the energy spectrum.
Given the students’ positive feedback regarding the course, particularly highlighting the exposure to various new topics such as cryogenics, light-matter interaction, and superconductivity, as well as the acquisition of skills in coding, data analysis, and problem-solving, the experiment has evolved into the primary student-led initiative within the new laboratory course at the cryogenics laboratories of the University of Milano-Bicocca. This trend has persisted throughout the years 2023 and 2024 and is anticipated to continue.
@@ Acknowledgments
Special thanks go to our laboratory supervisors for the opportunity and their guidance in our experimental activities. We acknowledge Benno Margesin and the FBK clean-room team for the chip fabrication. Additionally, we also acknowledge Renato Mezzena and Andrea Vinante for providing the device.
Data availability
The data that support the findings of this study are available upon request from the authors.
@@
[…]